Grade 11Mechanics

what are the equations kinematics explain kinematics

shaurya
6 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Hiii the kinematics euations are 
v = u + at 
Delta x = {(v+u)/2}*t
Delta X = ut + ½ at^2
v^2 = u^2 + 2as 
Kinematic formula are liisted on 5 different variables 
u – initial velocity 
s = distance 
v = final velocity 
a = acceleration 
t = time